OpenStudy (anonymous):

what is the y coordinate of the vertex of the parabola that is givenby the equation bellow: x = -2.5(x-3)^2 + 5

OpenStudy (mathmate):

The general equation in vertex form is given by: \((y-k)=a(x-h)^2\) for a parabola with vertical axis or \((x-h)=a(y-k)^2\) for a parabola with horizontal axis. where (h,k) represents the vertex. I suspect there is a typo in the question where the equation should read: \(y=-2.5(x-3)^2+5\) which is a parabola with vertical axis

OpenStudy (anonymous):

sorry I typed it wrong, its \[x = -2.5(y-3)^2 + 5\]

OpenStudy (mathmate):

In that case, it is a parabola with a horizontal axis. But the coordinates (h,k) still apply. It looks approximately like this.